find the equations of the foliowing lines in slope=intercept form the problem reads as this ; through (4,1) (1,3)
Answer by Alan3354(69443)

find the equations of the foliowing lines in slope=intercept form the problem reads as this ; through (4,1) (1,3)
------------------
Here's one similar.
write an equation in slope-intercept form of a line containing the points (-8,9)and (-6,-5).
--------------
There are 2 ways to do these.
Method 1:
Find the slope, m.
m = diffy/diffx
m = (-5-9))/(-6-(-8)) = -14/2
m = -7
Then use either point:
y-y1 = m*(x-x1)
y-9 = -7*(x+8)
y-9 = -7x-56
y = -7x-47
--------------
--------------
Method 2:
|+x +y +1|
|-8 +9 +1| = 0
|-6 -5 +1|
----------
x*(9+5) - y*(-8+6) + (40+54) = 0
14x + 2y +94 = 0
y = -7x - 47
--------------
Do it like that. If you need help or want me to check your answer email me via the Thank You note.